Hello,
How would I solve the following:
The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of its original thorium remaining. How old is the meteorite?
I think this has been answered but I can not find the answer.
Please help me I have tried several time to solve this without any help but I just can seen to solve it.
Thank you,
Jimmy

Answer by stanbon(75887)

The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of its original thorium remaining. How old is the meteorite?
-------------------------
A(t) = Ao(1/2)^(t/4.2 mil)
This says "the amount at time t A(t) is the amount you start with (Ao)
multiplied by (1/2) (t/4.2 mil) times.
-------------
0.076Ao = Ao*(1/2)^(t/4.2 mil)
0.076 = (1/2)^(t/4.2 mil)
Take the log of both sides to get:
log 0.076 = (t/4.2 mil)*log(1/2)
t/4.2 mil = log0.076/log0.5
t = 4.2 mil * 3.717857...
t = 15.614998.. million years old.
================
Cheers,
Stan H.